OpenStudy (anonymous):

Prove by mathematical induction, the following formula for the sum of a geometric series: S(n) r^i = r^(n+1) - 1/r-1 if r is not equal to 1

OpenStudy (perl):

first do the basis case,

OpenStudy (perl):

Sum r^i , i=1 .. n = r^(n+1) - 1/r-1

OpenStudy (anonymous):

\[\sum_{n}^{i=0} r^i = r ^{n-1} - 1 \div r-1,\] if \[r \neq 1\]

OpenStudy (anonymous):

Just to be clear. :-D

OpenStudy (anonymous):

did the sum upside down...I'm new here.

OpenStudy (anonymous):

How do I show the base case?

OpenStudy (perl):

ok plug in k=1

OpenStudy (perl):

also you should probably use parenthesees

OpenStudy (anonymous):

So, would the base case be n=0 or n=1?

OpenStudy (perl):

i believe i=0

OpenStudy (anonymous):

Yes, but in order to derive the base case, how would I format it?

OpenStudy (perl):

so start at i = 0

OpenStudy (perl):

\[\sum_{i=0}^{n-1} r^i = (r ^{n} - 1)/ (r-1)\ \]

OpenStudy (anonymous):

ok. So \[r ^{0}+r^{1} ... r^{n-1} ?\]

OpenStudy (perl):

the basis case is just one number

OpenStudy (perl):

lets do the basis case i = 1 then show r^0 + r^1 = ( r^2 - 1) / ( r - 1 )

OpenStudy (perl):

actually it might be easier if we do this

OpenStudy (perl):

\[\sum_{i=0}^{n} r^i = (r ^{n+1} - 1)/ (r-1)\ \]

OpenStudy (perl):

this is easier to prove, methinks

OpenStudy (perl):

with me so far?

OpenStudy (anonymous):

So n=1, \[r^{0} + r^{1} = (r^{(1+1)} - 1) / (r-1)\]

OpenStudy (perl):

so let's prove that r^0 + r^1 + r^2 + ... r^n = ( r^(n+1) - 1 ) ( r-1)

OpenStudy (perl):

we can use zero i = 0 r^0 = ( r^(1+0) - 1 ) / ( r - 1 )

OpenStudy (perl):

but i=1 is fine too, any number in fact

OpenStudy (anonymous):

Sorry to be obtuse, but what I'm confused about is that i is already = 0 as stated in the original equation

OpenStudy (anonymous):

So are you meaning to change n=0?

OpenStudy (perl):

'i' is the index, i = 0,1,2,3... n

OpenStudy (perl):

you dont have to apologize, any questions you can ask. just bear with me, this program is so buggy

OpenStudy (anonymous):

Yes. So in this case, it's stated that i=0. I thought in the base case you're supposed to set n to the lowest possible integer to test teh hypothesis

OpenStudy (anonymous):

So if we set n=0 we get: \[r^{0} = r^{1} - 1 / r-1\]

OpenStudy (anonymous):

which is 1=1. Would that be acceptable base case?

OpenStudy (perl):

yes,

OpenStudy (perl):

so it passes the basis case.

OpenStudy (perl):

sorry, the basis test

OpenStudy (anonymous):

Does the common ratio (r) not being equal to 1 factor into the basis test at all?

OpenStudy (perl):

if r = 1 , then we have division by zero

OpenStudy (anonymous):

Ok, but the basis test still holds true in this case.

OpenStudy (perl):

right

OpenStudy (perl):

because we assumed originally that r != 1

OpenStudy (anonymous):

so now to prove by induction, n = k+1 correct?

OpenStudy (anonymous):

or rather, we assume that \[\sum_{i=0}^{k} = r^{(k+1)}-1 / r-1\] is true

OpenStudy (perl):

right, and then show it will be true for \[\sum_{i=0}^{k+1} = (r^{((k+1)+1)}-1) / r-1\]

OpenStudy (anonymous):

and try by induction to prove that \[\sum_{i=0}^{k+1} r^{i} = r^{((k+1)+1)} -1 / r-1\]

OpenStudy (perl):

exactly

OpenStudy (anonymous):

Yayy I'm getting smarter!

OpenStudy (perl):

you just need parenthesees!

OpenStudy (anonymous):

So now, I have to somehow simplify it. \[\sum_{i=0}^{k+1} r^{i} = r^{0} + r ... r^{k+1} = (r^{k+2} - 1) / (r -1) \]

OpenStudy (perl):

sometimes its easier to write out the sum we assume that 1 + r + r^2 + ... + r^k = ( r^(k+1) - 1 ) / ( r - 1)

OpenStudy (perl):

what if we add r^(k+1) to both sides

OpenStudy (anonymous):

Oh wait, yes, I remember my professor doing that.

OpenStudy (perl):

then the left side is the left hand side of the conclusion

OpenStudy (perl):

and its a legal algebraic move

OpenStudy (anonymous):

Ok, can you explain that one more time? (Thank you so much for your time, this is exactly the clarification i needed)

OpenStudy (perl):

so i wrote out the assumption (what we assume to be true) 1 + r + r^2 + ... + r^k = ( r^(k+1) - 1 ) / ( r - 1) Then we stare at the conclusion 1 + r + r^2 + ... +r^k + r^(k+1) = ( r^(k+2) - 1 ) / ( r - 1) to go from the assumption to the conclusion, it seems natural to add r^k+1 to both sides of the assumption (and thats fine because we assumed it is true, and you can add the same number to both sides of a true equality)

OpenStudy (perl):

by adding r^k+1 to both sides of the assumption , that will immediately give us the LHS (left hand side) of the conclusion

OpenStudy (perl):

1 + r + r^2 + ... + r^k = ( r^(k+1) - 1 ) / ( r - 1) +r^k+1 +r^k+1 --------------------------------------------

OpenStudy (anonymous):

OK. so how can we simplify it from there to verify the equality?

OpenStudy (perl):

ok so we get

OpenStudy (perl):

1 + r + r^2 + ... + r^k = ( r^(k+1) - 1 ) / ( r - 1) +r^k+1 +r^k+1 -------------------------------------------- 1+ r + r^2 + ... + r^k +1 = ( r^(k+1) - 1 ) / ( r - 1) + r^k+1

OpenStudy (perl):

now we just have to simplify the right hand side (RHS)

OpenStudy (perl):

did i lose you?

OpenStudy (anonymous):

No, I'm with you so far.

OpenStudy (anonymous):

Actually

OpenStudy (anonymous):

in the RHS why is the parenthetical \[r^{k+1}\] not \[r^{k+2}\]

OpenStudy (perl):

because that was what we had in the assumption

OpenStudy (perl):

but after we simplify we should get r^k+2

OpenStudy (anonymous):

Ok so how do we simplify the RHS

OpenStudy (anonymous):

because this is where I always screw up

OpenStudy (perl):

we just need a common denominator

OpenStudy (perl):